OpenStudy (anonymous):

What is the solution set of |2x – 3| – 6 = –1?

myininaya (myininaya):

Add 6 to both sides first.

myininaya (myininaya):

|2x-3|=5

myininaya (myininaya):

=> 2x-3=5 or 2x-3=-5

myininaya (myininaya):

Solve both of these equations and you will have your answers

OpenStudy (vishal_kothari):

-1,4...

OpenStudy (anonymous):

What is the solution set of |2x – 3| – 6 = –1? the | | signs mean just the simple answer, 2x-3, if it had a negative sign in front loike |-2x-3| it will still be 2x-3 the answer being used to solve! so dont let that confuse you!! ok first take out the || sign. so 2x-3-6= -1 ... now solve! 2x-9=-1...... add 9 to both sides so, it will cancel out in the left and ur left with 2x=8 now divide by 2 in both sides and ur left with x=4!!!!

OpenStudy (anonymous):

2x-3 is a absolute value from line ,so it has both answer .

myininaya (myininaya):

|2x-3|=2x-3 when x>3/2 |2x-3|=-(2x-3) when x<3/2 we must look at both cases |2x-3|-6=-1 |2x-3|=6-1 |2x-3|=5 => 2x-3=5 or 2x-3=-5 => 2x=8 or 2x=-2 => x=4 or x=-1 Check both, kmila, these are both solutions!

myininaya (myininaya):

check x=-1 we have |2(-1)-3|-6=|-2-3|-6=|-5|-6=5-6=-1 <- we have this on both sides of equation => x=-1 is a solution check x=4 |2(4)-3|-6=|8-3|-6=|5|-6=5-6=-1 <- we have this on both sides of equation => x=4 is a solution => we have two solutions x={-1,4}